\begin{thebibliography}{AI}

\bibitem{Pu52} Pao Ming Pu, Some inequalities in certain nonorientable
Riemannian manifolds, Pacific J. Math.  2 (1952), 55--71.


\bibitem{Gr83} Mikhael Gromov, Filling Riemannian manifolds.
J. Differential Geom. 18 (1983), no.\;1, 1--147.


\bibitem{Gr99} Mikhael Gromov, Metric structures for Riemannian and
non-Riemannian spaces, Based on the 1981 French original with
appendices by M. Katz, P. Pansu and S. Semmes, Translated from the
French by Sean Michael Bates, Progress in Mathematics, 152,
Birkh\"auser Boston, Boston, MA, 1999.


\bibitem{Be65} Marcel Berger, Lectures on geodesics in Riemannian
geometry, Tata Institute of Fundamental Research Lectures on
Mathematics, No. 33 Tata Institute of Fundamental Research, Bombay
1965.


\bibitem{Ho09} Charles Horowitz, Karin Usadi Katz, Mikhail Katz,
Loewner's torus inequality with isosystolic defect, J. Geom. Anal.  19
(2009), no.\;4, 796--808.



\bibitem{Ba19} Ivan Babenko, Florent Balacheff, and Guillaume Bulteau,
Systolic geometry and simplicial complexity for groups, J. Reine
Angew. Math.  757 (2019), 247--277.


\bibitem{Ba10} Florent Balacheff, A local optimal diastolic inequality
on the two-sphere, J. Topol. Anal. 2 (2010), no.\;1, 109--121.

\bibitem{Ba07} Viktor Bangert, Christopher Croke, Sergei Ivanov, and
Mikhail Katz, Boundary case of equality in optimal Loewner-type
inequalities, Transactions of the American Mathematical Society
\textbf{359} (2007) (1), 1--17.



\bibitem{Cr20} Paul Creutz,   Rigidity of the Pu inequality and quadratic
isoperimetric constants of normed spaces, preprint (2020).  See
\url{https://arxiv.org/pdf/2004.01076.pdf}




\bibitem{El15} Chady El Mir and Zeina Yassine, Conformal geometric
inequalities on the Klein bottle, Conform. Geom. Dyn. 19 (2015),
240--257.



\bibitem{Gu10} Larry Guth, Systolic inequalities and minimal
hypersurfaces, Geom. Funct. Anal.  19 (2010), no. 6, 1688--1692.



\bibitem{He19} James Hebda, The primitive length spectrum of 2-D tori
and generalized Loewner inequalities, Trans. Amer. Math. Soc. 372
(2019), no.\;9, 6371--6401.


\bibitem{Iv04} Sergei Ivanov and Mikhail Katz, Generalized degree and
optimal Loewner-type inequalities, Israel J. Math.  141 (2004),
221--233.


\bibitem{Ka07} Mikhail Katz, Systolic geometry and topology, with an
appendix by Jake P. Solomon, Mathematical Surveys and Monographs, 137,
American Mathematical Society, Providence, RI, 2007.


\bibitem{Ka06} Mikhail Katz and Stephane Sabourau, Hyperelliptic
surfaces are Loewner, Proceedings of the American Mathematical Society
134 (2006), no.\;4, 1189--1195.


\bibitem{Li17} Yevgeny Liokumovich, Alexander Nabutovsky, and Regina
Rotman, Lengths of three simple periodic geodesics on a Riemannian
2-sphere, Math. Ann.  367 (2017), no. 1--2, 831--855.


\bibitem{Sa11} Stephane Sabourau, Isosystolic genus three surfaces
critical for slow metric variations, Geom. Topol.  15 (2011), no.\;3,
1477--1508.


\bibitem{Cs18} Bal\'azs Csik\'os and M\'arton Horv\'ath, Harmonic
manifolds and tubes, J. Geom. Anal. 28 (2018), no.\;4, 3458--3476.



\end{thebibliography}

\documentclass[12pt,reqno]{amsart}

\usepackage[arrow,matrix,curve]{xy}

\usepackage[dvips]{graphicx} 

\usepackage{amssymb, latexsym, amsmath, amscd, array, hyperref
}

\usepackage{breakurl}   

\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{corollary}[theorem]{Corollary}

\theoremstyle{definition}

\numberwithin{equation}{section}

\DeclareMathOperator{\area}{{\rm area}}

\DeclareMathOperator{\sys}{{\rm sys}}
\DeclareMathOperator{\Var}{{\rm Var}}

\newcommand\R {{\mathbb R}}

\newcommand{\RP}{{\mathbb R\mathbb P}}

\numberwithin{equation}{section}

\author{Mikhail G. Katz}\address{M. Katz, Department of Mathematics,
Bar Ilan University, Ramat Gan 5290002
Israel}\email{katzmik@macs.biu.ac.il}

\author{Tahl Nowik}\address{T. Nowik, Department of Mathematics, Bar
Ilan University, Ramat Gan 5290002 Israel}\email{tahl@math.biu.ac.il}

\title [Systolic inequality with remainder in projective plane] {A
systolic inequality with remainder in the real projective plane}

\begin{document}

\thispagestyle{empty}


\begin{abstract}
The first paper in systolic geometry was published by Loewner's
student P. M. Pu over half a century ago.  Pu proved an inequality
relating the systole and the area of an arbitrary metric on the real
projective plane.  We prove a stronger version of Pu's systolic
inequality with a remainder term.
\end{abstract}

\keywords{Systole; geometric inequality; Riemannian submersion;
Cauchy-Schwarz theorem; probabilistic variance}

\subjclass[2010]{Primary 53C23; Secondary 
53A30 
}

\maketitle


\section{Introduction}

Loewner's systolic inequality for the torus and Pu's inequality
\cite{Pu52} for the real projective plane were historically the first
results in systolic geometry.  Great stimulus was provided in 1983 by
Gromov's paper \cite{Gr83}, and later by his book \cite{Gr99}.

Our goal is to prove a strengthened version with a remainder term of
Pu's systolic inequality~$\sys^2(g)\leq\frac\pi2\area(g)$ (for an
arbitrary metric~$g$ on~$\RP^2$), analogous to Bonnesen's
inequality~$L^2 - 4\pi A \geq \pi^2(R-r)^2$, where~$L$ is the length
of a Jordan curve in the plane,~$A$ is the area of the region bounded
by the curve,~$R$ is the circumradius and~$r$ is the inradius.

Note that both the original proof in Pu (\cite{Pu52}, 1952) and the
one given by Berger (\cite{Be65}, 1965, pp.\;299--305) proceed by
averaging the metric and showing that the averaging process decreases
the area and increases the systole.  Such an approach involves a
5-dimensional integration (instead of a 3-dimensional one given here)
and makes it harder to obtain an explicit expression for a remainder
term.  Analogous results for the torus were obtained in \cite{Ho09}
with generalisations in \cite{Ba19} - \cite{Sa11}.

\section{The results}

We define a closed~$3$-dimensional manifold~$M \subseteq \R^3 \times
\R^3$ by setting
\[
M=\{ (v,w) \in \R^3 \times \R^3 \colon \; v\cdot v =1, \ w\cdot w=1, \
v\cdot w=0 \}
\]
where~$v \, \cdot \, w$ is the scalar product on~$\R^3$. We have a
diffeomorphism~$M\to SO(3,\R)$,~$(v,w) \mapsto (v,w,v\times w)$,
where~$v\times w$ is the vector product on~$\R^3$.  Given a
point~$(v,w) \in M$, the tangent space~$T_{(v,w)}M$ can be identified
by differentiating the three defining equations of~$M$ along a path
through~$(v,w)$.  Thus
\[
T_{(v,w)}M=\{ (X,Y) \in \R^3 \times \R^3\colon X\cdot v =0, \ Y\cdot
w=0, \ X\cdot w+Y\cdot v=0 \}.
\]

We define a Riemannian metric~$g_M$ on~$M$ as follows.  Given a point
$(v,w)\in M$, let~$n=v\times w$ and declare the basis~$(0,n) , (n,0) ,
(w,-v)$ of~$T_{(v,w)}M$ to be orthonormal.  This metric is a
modification of the metric restricted to~$M$
from~$\R^3\times\R^3=\R^6$.  Namely, with respect to the Euclidean
metric on~${\R}^6$ the above three vectors are orthogonal and the
first two have length 1.  However, the third vector has Euclidean
length~$\sqrt{2}$, whereas we have defined its length to be 1.  Thus
if~$A\subseteq T_{(v,w)}M$ denotes the span of~$(0,n)$ and~$(n,0)$,
and~$B\subseteq T_{(v,w)}M$ is spanned by~$(w,-v)$, then the metric
$g_M$ on~$M$ is obtained from the Euclidean metric~$g$ on~$\R^6$
(viewed as a quadratic form) as follows:
\begin{equation}
g_M = g\!\downharpoonright_A^{\phantom{I}}+\,\frac12\,
g\!\downharpoonright_B^{\phantom{II}}.
\end{equation}

Each of the natural projections~$p,q \colon M \to S^2$ given
by~$p(v,w)=v$ and~$q(v,w)=w$, exhibits~$M$ as a circle bundle
over~$S^2$.

\begin{lemma}
\label{l101}
The maps~$p$ and~$q$ on~$(M,g_M)$ are Riemannian submersions, over the
unit sphere~$S^2\subseteq\R^3$.
\end{lemma}

\begin{proof}
For the projection~$p$, given~$(v,w) \in M$, the vector~$(0,n)$ as
defined above is tangent to the fiber~$p^{-1}(v)$. The other two
vectors,~$(n,0)$ and~$(w,-v)$, are thus an orthonormal basis for the
subspace of~$T_{(v,w)}M$ normal to the fiber, and are mapped by~$dp$
to the orthonormal basis~$n,w$ of~$T_v S^2$.
\end{proof}

The projection~$p$ maps the fiber~$q^{-1}(w)$ onto a great circle
of~$S^2$.  This map preserves length since the unit vector~$(n,0)$,
tangent to the fiber~$q^{-1}(w)$ at~$(v,w)$, is mapped by~$dp$ to the
unit vector~$n \in T_v S^2$.  The same comments apply when the roles
of~$p$ and~$q$ are reversed.

In the following proposition, integration takes place respectively
over great circles~$C\subseteq S^2$, over the fibers in~$M$,
over~$S^2$, and over~$M$.  The integration is always with respect to
the volume element of the given Riemannian metric.  Since~$p$ and~$q$
are Riemannian submersions by Lemma~\ref{l101}, we can use Fubini's
Theorem to integrate over~$M$ by integrating first over the fibers of
either~$p$ or~$q$, and then over~$S^2$; cf.\;\cite[Lemma\;4]{Cs18}.
By the remarks above, if~$C=p(q^{-1}(w))$ and~$f\colon S^2 \to \R$
then~$\int_{q^{-1}(w)} f\circ p = \int_C f$.

\begin{proposition}
\label{inq}
Given a continuous function~$f\colon S^2 \to \R^+$, we define~$m\in\R$
by setting~$m=\min \{\int_C f\colon C \subseteq S^2 \text{ a great
circle} \}$.  Then
\[
\frac{m^2}{\pi}\leq\frac{1}{4\pi}\bigg(\int_{S^2}f\bigg)^2\leq\int_{S^2}f^2,
\]
where equality in the second inequality occurs if and only if~$f$ is
constant.
\end{proposition}

\begin{proof}
Using the fact that~$M$ is the total space of a pair of Riemannian
submersions, we obtain
\[
\begin{aligned}
\int_{S^2} f & = \int_{S^2}\bigg(\frac{1}{2\pi}\int_{p^{-1}(v)} f\circ
p\bigg)
\\&= \frac{1}{2\pi}\int_M f \circ p 
\\&=
\frac{1}{2\pi}\int_{S^2}\bigg(\int_{q^{-1}(w)} f\circ p\bigg) \\&\geq
\frac{1}{2\pi}\int_{S^2}m =2m,
\end{aligned}
\]
proving the first inequality.  By the Cauchy--Schwarz inequality, we
have
\[
\Big(\int_{S^2} 1 \cdot f \Big)^2 \leq 4 \pi \int_{S^2} f^2,
\]
proving the second inequality.  Here equality occurs if and only if
$f$ and~$1$ are linearly dependent, i.e., if and only if~$f$ is
constant.
\end{proof}

We define the quantity~$V_f$ by setting~$V_f =\int_{S^2} f^2 -
\frac{1}{4\pi}\big(\int_{S^2} f \big)^2$.  Then Proposition \ref{inq}
can be restated as follows.

\begin{corollary}
\label{c1}
Let~$f\colon S^2 \to \R^+$ be continuous.  Then
\[
\int_{S^2} f^2 - \frac{m^2}{\pi} \geq V_f \geq 0,
\]
and~$V_f=0$ if and only if~$f$ is constant.
\end{corollary}

\begin{proof}
The proof is obtained from Proposition~\ref{inq} by noting that~$a\leq
b \leq c$ if and only if~$c-a \geq c-b \geq 0$.
\end{proof}

We can assign a probabilistic meaning to~$V_f$ as follows. Divide the
area measure on~$S^2$ by~$4\pi$, thus turning it into a probability
measure~$\mu$.  A function~$f\colon S^2 \to \R^+$ is then thought of
as a random variable with
expectation~$E_\mu(f)=\frac{1}{4\pi}\int_{S^2}f$.  Its variance is
thus given by
\[
\Var_\mu(f)=E_\mu(f^2)-\big(E_\mu(f)\big)^2 = \frac{1}{4\pi}\int_{S^2}
f^2 - \bigg(\frac{1}{4\pi}\int_{S^2} f\bigg)^2 = \frac{1}{4\pi} V_f.
\]
The variance of a random variable~$f$ is non-negative, and it vanishes
if and only if~$f$ is constant. This reproves the corresponding
properties of~$V_f$ established above via the Cauchy--Schwarz
inequality.

Now let~$g_0$ be the metric of constant Gaussian curvature~$K=1$
on~$\RP^2$.  The double covering~$\rho \colon S^2 \to (\RP^2, g_0)$ is
a local isometry.  Each projective line~$C\subseteq\RP^2$ is the image
under~$\rho$ of a great circle of~$S^2$.

\begin{proposition}
\label{inqP}
Given a function~$f\colon \RP^2 \to \R^+$, we define~$\bar m\in\R$ by
setting~$\bar{m}=\min \{ \int_C f \colon C \subseteq \RP^2 \ \text{a
projective line} \}$.  Then
\[
\frac{2\bar{m}^2}{\pi} \ \leq \ \frac{1}{2\pi}\bigg(\int_{\RP^2} f
\bigg)^2 \ \leq \ \int_{\RP^2} f^2,
\]
where equality in the second inequality occurs if and only if~$f$ is
constant.
\end{proposition}

\begin{proof}
We apply Proposition \ref{inq} to the composition~$f\circ\rho$.  Note
that we have $\int_{\rho^{-1}(C)} f\circ \rho = 2 \int_C f$
and~$\int_{S^2} f \circ \rho = 2 \int_{\RP^2} f$. The condition
for~$f$ to be constant holds since~$f$ is constant if and only
if~$f\circ \rho$ is constant.
\end{proof}

For~$\RP^2$ we define~$\bar{V}_f = \int_{\RP^2} f^2 -
\frac{1}{2\pi}\big(\int_{\RP^2} f \big)^2 =\frac{1}{2}V_{f\circ\rho}$.
We obtain the following restatement of Proposition \ref{inqP}.

\begin{corollary}
\label{c1P} 
Let~$f\colon\RP^2 \to \R^+$ be a continuous function.  Then
\[
\int_{\RP^2} f^2 - \frac{2\bar{m}^2}{\pi} \geq \bar{V}_f \geq 0,
\]
where~$\bar{V}_f=0$ if and only if~$f$ is constant.
\end{corollary}

Relative to the probability measure induced by~$\frac{1}{2\pi}g_0$
on~$\RP^2$, we have~$E(f) = \frac{1}{2\pi}\int_{\RP^2}f$, and
therefore~$\Var(f)=\frac{1}{2\pi}\bar{V}_f$, providing a probabilistic
meaning for the quantity~$\bar{V}_f$, as before.

By the uniformization theorem, every metric~$g$ on~$\RP^2$ is of the
form~$g=f^2 g_0$ where~$g_0$ is of constant Gaussian curvature~$+1$,
and the function~$f\colon\RP^2\to\R^+$ is continuous.  The area of~$g$
is~$\int_{\RP^2} f^2$, and the~$g$-length of a projective line~$C$ is
$\int_C f$.  Let~$L$ be the shortest length of a noncontractible
loop. Then~$L\leq \bar{m}$ where~$\bar{m}$ is defined in
Proposition~\ref{inqP}, since a projective line in~$\RP^2$ is a
noncontractible loop.  Then Corollary~\ref{c1P} implies~$\area(\RP^2,
g) - \frac{2L^2}{\pi} \ \geq \ \bar{V}_f \ \geq \ 0$.
If~$\area(\RP^2, g) = \frac{2L^2}{\pi}$ then~$ \bar{V}_f = 0$, which
implies that~$f$ is constant, by Corollary \ref{c1P}. Conversely, if
$f$ is a constant~$c$, then the only geodesics are the projective
lines, and therefore~$L=c\pi$.  Hence~$\frac{2L^2}{\pi} = 2\pi
c^2=\area(\RP^2)$.  We have thus completed the proof of the following
result strengthening Pu's inequality.

\begin{theorem}
\label{pu}
Let~$g$ be a Riemannian metric on~$\RP^2$.  Let~$L$ be the shortest
length of a noncontractible loop in~$(\RP^2,g)$.
Let~$f\colon\RP^2\to\R^+$ be such that~$g=f^2g_0$ where~$g_0$ is of
constant Gaussian curvature~$+1$.  Then 
\[
\area(g) - \frac{2L^2}{\pi}\geq 2\pi \Var(f),
\]
where the variance is with respect to the probability measure induced
by~$\frac{1}{2\pi} g_0$.  Furthermore,
equality\,~$\area(g)=\frac{2L^2}{\pi}$ holds if and only if~$f$ is
constant.
\end{theorem}

\bibliographystyle{amsalpha}
 
\end{document}